A group of MRU students starts working on a research project about lowering the speed limit on Calgary’s residential streets to 30 km/h. They record the speed of the traffic on Richard Road SW with a 30 km/h speed limit. The following table represents their results.

Speed of Car

Frequency

10-15

2

16-21

4

22-27

12

28-33

29

34-39

22

40-45

18

46-51

16

52-57

10

58-63

7

  1. Identify the lower class limit for each class.

  1. Identify the upper class limit for each class.

  1. What is the class width?

  1. How many cars were involved in this research project?

  1. What class does represent the most common speed range?

  1. Find the percentage of the cars having a speed between 52 and 57 km/h.
Answer #1
lower class upper class
10 15
16 21
22 27
28 33
34 39
40 45
46 51
52 57
58 63

class width = upper class - lower class = 15-10 = 5

How many cars were involved in this research project?

sum of frequency = 120

The class 28-33 is represent most common speed of class = 29

Find the percentage of the cars having a speed between 52 and 57 km/h. = 10/120 = 0.0833

= 8.33%